OpenStudy (anonymous):

Solve. 36x+9x^2=0. I came up with 0,-4. Can someone confirm.

OpenStudy (anonymous):

that's correct. (:

OpenStudy (anonymous):

ok. thanks. medal

OpenStudy (anonymous):

9(x^2+4x) =0 9(x(x+4)=0 x+4 = 0 x =-4 x =0

OpenStudy (anonymous):

Thank

OpenStudy (anonymous):

good cluck your home work